OpenStudy (anonymous):

Find a third-degree polynomial equation with rational coefficients that has roots -4 and 6+i. When I solved this, I got the answer x^3-8x^2-13x+140. This is not one of the answers. Can someone tell me what I did wrong?

OpenStudy (anonymous):

*fanning me before I even help? haha ^_^

OpenStudy (anonymous):

You helped me on a previous problem yesterday and I forgot to fan you because I couldn't find you! Haha

OpenStudy (anonymous):

oh did i? lol

OpenStudy (anonymous):

I know you started by doing (x+4)(x-(6+i))(x-(6-1))

OpenStudy (anonymous):

And then once you get passed all of that... you get (x+4)(x^2-12x+35)

OpenStudy (anonymous):

jk, ur right

OpenStudy (anonymous):

What am I right about? Lol

OpenStudy (anonymous):

give me a sec..

OpenStudy (anonymous):

Okay! Thank you

OpenStudy (anonymous):

"has roots -4 and 6+i." doesn't this mean these are the only possible roots?

OpenStudy (anonymous):

I'll send you the possible answer choices

OpenStudy (anonymous):

x^3-8x-11x+148=0 x^3-8x^2-12x+37=0 x^3-12x+37=0 x^2-11x=0

OpenStudy (anonymous):

Oops, the last one is x^3-8x^2-11x=0

OpenStudy (anonymous):

I got what you got... O.o

OpenStudy (anonymous):

Hmmm...

OpenStudy (anonymous):

I don't know what I did wrong...

OpenStudy (anonymous):

when you foil your complex conjugates you shouldn't have a middle term with an x

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

well yyou would get (x+4)(x^2-12x+37) then foil that

OpenStudy (anonymous):

remember -I^2=+1

OpenStudy (anonymous):

you get x^3-8x^2-11x+148

OpenStudy (anonymous):

That's what I got too! Okay. Thank you!!

OpenStudy (anonymous):

no problem. little 1 number mistakes can mess with you.

OpenStudy (anonymous):

Yes they can... I figured it was something like that